Session: Nov 1998 (session 4)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (M. Bondi at Fri Nov 13 05:17:38 MET 1998) Success
No problems in running the observation

Torun (K.Borkowski at Fri Nov 13 09:37:53 MET 1998) Minor failures
1730 to 1747 no useful observations due to a loss of the communication with the formatter. Also, the schedule did not allow for slewing time while changing the observed sources.

Onsala (A. Polatidis at Fri Nov 13 15:59:15 MET 1998) Success
Intermittent interference across the band, but it probably does not affect the data too much. The interference lasts for less than a second each time, it is probably a radar. Lost a few seconds while the telescope was slewing from source to source. The tape will be sent to Socorro on Tuesday.

Effelsberg (dgraham at Mon Nov 16 16:38:29 MET 1998) Minor failures
no problems noticed

Jodrell (Alastair Gunn at Wed Nov 18 11:24:15 MET 1998) Minor failures
Cable wrap at approx 01:00 UT, telescope on source again at approx 01:17 UT. No other known problems.

Calibration (Jean\-Francois Desmurs at Tue Nov 24 19:05:43 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: ed010.antab.Read the comments in the file to get further information.